OpenStudy (anonymous):

Let f(x) = 7x - 13. Find f-1(x)

OpenStudy (anonymous):

the 1 is an exponent

OpenStudy (anonymous):

f^(-1) (x) is the inverse function of x. To find the inverse function, interchange x with y, and y with x, then solve for y. In your problem, y = 7x - 13 So the inverse is x = 7y - 13. Now solve for y, and that will be the inverse function of y = 7x - 13.

OpenStudy (anonymous):

what's your inverse function?

OpenStudy (anonymous):

JAMALB..whats your answer?

OpenStudy (jack1):

sorry... what's the inverse function again...?

OpenStudy (anonymous):

JAMALB, yes, the mathematical notation for the inverse of a function f, is f^(-1) which the -1 is the exponent.

OpenStudy (jack1):

no, just up above @Easyaspi314 u said: In your problem, y = 7x - 13 So the inverse is x = 7y - 13.... this isn't the inverse

OpenStudy (anonymous):

I said that you need to solve for y in the equation x = 7y - 13 Read above what I wrote!

OpenStudy (anonymous):

I wanted JAMALB to solve that for y.

OpenStudy (anonymous):

im working it now

OpenStudy (anonymous):

ok

OpenStudy (jack1):

ah, my bad, i thought u meant that was ur inverse equation, all good then

OpenStudy (anonymous):

Best is that if you come in the middle of a conversation, first read all of the posts, then feel free to jump in.

OpenStudy (jack1):

no, it wasnt that, it was the phrasing of "So the inverse is x = 7y - 13" i did read from the top down

OpenStudy (anonymous):

x=7y-13 7y=x-13 y=x-13/7

OpenStudy (anonymous):

JAMALB, thats incorrect. x = 7y - 13 First, add 13 to both sides of the equation, you get x + 13 = 7y Now divide both sides by 7, you get y = (x + 13)/7

OpenStudy (ranga):

If you are interested in knowing the meaning of inverse function of f(x) = 7x - 13: In f(x) = 7x - 13, we can put various values for the independent variable x and compute f(x). With inverse function we are asking if we are given f(x) how will we compute x?